In a shipment of 20 computers, 2 are defective. Three computers are randomly selected and tested. What is the probability that all 2 are defective if the first one is not replaced after being tested?
1
Expert's answer
2020-12-15T07:04:56-0500

Probabilities of being defective in case the 1st is not replaced (2nd is replaced)

First one: "\frac{2}{20}"

Second one: "\frac{1}{19}"

Third one: no more defective


Therefore we have: "\frac{2}{20}\cdot\frac{1}{19}=\frac{1}{190}"